OpenStudy (anonymous):

Julie drives 100 mi to Grandmother's house. On the way to Grandmother's, Julie drives half the distance at 33.0 and half the distance at 80.0 . On her return trip, she drives half the time at 33.0 and half the time at 80.0 .

OpenStudy (anonymous):

What is Julie's average speed on the way to Grandmother's house?

OpenStudy (anonymous):

What total distance did she travel?

OpenStudy (turingtest):

(total distance)/(total time)=average speed each travel time is distance over velocity t1=50/33 t2=50/80 so then we plug in the numbers into the formula average speed =100/(t1+t2)